A corporation pays an annual dividend of 10% on its preferred stock whose face value is $50. If similar companies pay a return of 15%, how much is a share of the corporation's preferred stock worth?
Answer #1

Calculation of price of preferred stock:

Dividend= 50*0.10= 5

Required return= 15%

Current price= dividend/required return

Current price= 5/0.15= 33.33

Price of preferred stock is 33.33